# Language Model Preference Evaluation with Multiple Weak Evaluators

### v2

@@ A.5 Definition of Cycle Rate
A.6 Dataset
@@
A.7 Aggregation Process in GED Across Different Tasks
@@ Appendix B Alternative Evaluator Configurations for Model Ranking
Appendix C Impact of Evaluator Quantity on Denoising Quality
@@ Appendix D Construction of the Preference Graph
1:Set of candidate answers $\{ans_{1},ans_{2},\ldots,ans_{n}\}$ for each question $q\in Q$, set of evaluators $\mathcal{A}=\{a_{1},a_{2},\ldots,a_{k}\}$
3:for each question $q\in Q$ do
4:     for each evaluator $a\in\mathcal{A}$ do
5:         Initialize vertex set $V_{q}=\{v_{1},v_{2},\ldots,v_{n}\}$, where each $v_{i}$ corresponds to $ans_{i}$
7:         for each pair of answers $(ans_{i},ans_{j})$ with $i\neq j$ do
8:              if $a(ans_{i},ans_{j})>0$ then
9:                  if $(v_{i},v_{j})\notin A_{a}$ then
10:                       Add directed edge $(v_{i}\rightarrow v_{j})$ to $A_{a}$, set $w_{a}(v_{i},v_{j})=1$
11:                  else
12:                       Increment $w_{a}(v_{i},v_{j})$ by 1
13:                  end if
14:              else
15:                  if $(v_{j},v_{i})\notin A_{a}$ then
16:                       Add directed edge $(v_{j}\rightarrow v_{i})$ to $A_{a}$, set $w_{a}(v_{j},v_{i})=1$
17:                  else
18:                       Increment $w_{a}(v_{j},v_{i})$ by 1
19:                  end if
20:              end if
21:         end for
22:         Store the preference graph $G_{a}=(V_{q},A_{a},w_{a})$ for evaluator $a$
23:     end for
24:end for
Appendix E Proof of Theorem 1
@@
We will then use contradiction to show it is impossible for $\mathrm{MAS}(G)$ to contain neither $(u\rightarrow v)$ nor $(v\rightarrow u)$. Suppose this is true instead. Then, since $\mathrm{MAS}(G)$ is a maximum acyclic subgraph, adding either $(u\rightarrow v)$ or $(v\rightarrow u)$ to $\mathrm{MAS}(G)$ will make it cyclic. That is, $\mathrm{MAS}(G)$ contains a path that goes from $v$ to $u$; meanwhile, it also contains a path from $u$ to $v$. As a result, it contains a cycle that goes from $v$ to $u$ and then from $u$ to $v$, which contradicts with the fact that $\mathrm{MAS}(G)$ is a maximum acyclic subgraph. Therefore, $\mathrm{MAS}(G)$ must contain exactly one of $(u\rightarrow v)$ and $(v\rightarrow u)$.
Appendix F Impact of Evaluator Weighting on GED Performance
@@
Appendix G Evaluating GED on Diverse Metrics
To address the need for more comprehensive assessments beyond accuracy-based metrics, we expanded our evaluation of GED to include nuanced quality aspects of LLM-generated outputs, such as factuality, relevance, coherence, informativeness, helpfulness, and validity. This evaluation was motivated by the understanding that tasks involving LLMs often require subtle judgments beyond simple accuracy, and GED’s adaptability to these scenarios is crucial. Building upon this foundation, we conducted experiments following (Chen et al., 2023a). Specifically, we used Llama3-70B to generate ten candidate responses for each query, and GED was applied in the response selection setting with Llama3-8B, Mistral-7B, and Qwen2-7B serving as evaluators. Each metric—factuality, relevance, coherence, informativeness, helpfulness, and validity—was assessed to measure GED’s ability to enhance the overall quality of LLM outputs. The results, shown in Table 7, highlight GED’s effectiveness. GED consistently outperformed individual evaluators and random selection across all metrics. For instance, GED improved factuality by approximately 5 percentage points over the best individual evaluator. Similarly, it enhanced relevance and coherence, indicating better alignment with the query and logically consistent responses. GED also demonstrated notable improvements in informativeness and helpfulness, suggesting that the responses selected by GED provide more valuable and user-centric information. These findings confirm GED’s adaptability to tasks requiring nuanced quality assessments. By aggregating preferences from multiple evaluators, GED captures subtle qualities of generated content, enhancing not only accuracy but also the overall reliability and utility of LLM outputs. This adaptability reinforces GED’s value in scenarios demanding comprehensive evaluations of language models.
Table 7: Performance comparison on nuanced quality metrics (in %). GED outperforms individual evaluators and random selection across Factuality, Relevance, Coherence, Inform., Helpful. and Validity metrics.
Random 86.73 87.91 92.47 77.62 17.48 48.92 68.52
Llama3-8B 88.59 89.91 94.41 79.77 18.48 50.92 70.35
Mistral-7B 89.10 90.29 94.85 79.95 18.55 51.13 70.65
Qwen2-7B 89.25 90.44 95.03 80.09 18.58 51.21 70.77
GED 94.73 95.91 97.36 86.62 19.48 55.92 75.00
Appendix H Expanded Related Work on Preference Denoising for LLMs
This work situates itself within the broader field of denoising preference evaluations for LLMs, addressing inconsistencies and noise in preference graphs. We acknowledge that existing literature has explored two primary approaches to this challenge: within-model preference modeling and modular pre-processing. Below, we provide a detailed comparison of GED with representative methods from these approaches.
@@ Appendix H Expanded Related Work on Preference Denoising for LLMs
Within-model approaches, such as robust DPO (rDPO) (Chowdhury et al., 2024) and conservative DPO (cDPO) (Mitchell, 2023), focus on integrating denoising mechanisms directly within the preference modeling process. These methods incorporate regularization techniques to mitigate the effects of noisy or adversarial preference data during model alignment. While these approaches are effective in refining the preference modeling pipeline, they are tightly coupled with specific models and tasks, limiting their versatility. In contrast, GED is a modular framework that operates as a pre-processing step, denoising preference graphs before downstream tasks, making it adaptable to a wide range of applications and models. Modular approaches like CURATRON (Naresh et al., 2024) address noise and missing comparisons in preference datasets using techniques such as low-rank matrix completion. While CURATRON effectively mitigates certain types of noise, it does not explicitly target cyclic inconsistencies (e.g., A $\succ$ B, B $\succ$ C, C $\succ$ A), which are a critical focus of GED. By leveraging a graph-based framework, GED detects and removes such cycles through its denoising process, ensuring that the resulting preference graph is acyclic and thus more consistent and reliable for downstream use. Additionally, GED distinguishes itself by providing theoretical guarantees for recovering the ground truth preference DAG under certain conditions. Furthermore, the ensemble mechanism in GED demonstrates the novel insight that combining weaker evaluators can surpass the performance of a single stronger evaluator, a feature not emphasized in the aforementioned methods.
Appendix I Impact of Evaluator Selection on GED Performance
@@
The selection of weak evaluators is a critical factor influencing the performance of GED. While our initial experiments demonstrated the benefits of combining weaker evaluators, further analysis is necessary to understand how their diversity and capabilities affect GED’s effectiveness. To address this, we conducted additional experiments evaluating three different configurations of evaluator sets: the original set comprising Llama3-8B, Mistral-7B, and Qwen2-7B; a diverse set that adds Gemma-9B [^4] https://huggingface.co/google/gemma-2-9b-it to introduce more variation in model architecture and training data; and a higher-capability set that replaces smaller models with larger ones, specifically Llama3-70B, Mistral-8×7B, and Qwen2-72B. The results, shown in Table 8, reveal two key insights. First, incorporating diversity by adding Gemma-9B improves GED’s performance slightly, increasing the average score from 56.05% to 56.39%. This suggests that models with diverse training data and architectures can contribute to more robust aggregated evaluations. Second, replacing smaller evaluators with higher-capability models yields a more substantial improvement, with the average score rising to 60.89%. Notably, benchmarks like AlpacaEval and GAIA benefit the most from the advanced reasoning and language understanding capabilities of larger models. These findings demonstrate the importance of both diversity and capability in evaluator selection. Diversity provides marginal gains by bringing varied perspectives, while higher-capacity models contribute to more significant improvements by enhancing the overall quality of evaluations. This suggests that, when computational resources allow, incorporating advanced models into the evaluator set can meaningfully boost GED’s performance.
@@
Appendix J Cost Considerations and Active Learning with ActiveGED
Evaluators Set of GED HumanEval AlpacaEval MATH GSM8k GAIA Avg.
Random 57.93 29.58 72.75 84.67 11.52 51.29
ActiveGED (30%) 67.28 30.96 74.65 85.73 11.39 54.00
ActiveGED (50%) 68.62 31.91 74.87 87.06 12.08 54.91
GED (Full Budget) 70.73 32.44 75.58 88.18 13.33 56.05
@@
Appendix K Mitigating Evaluator Biases in GED
@@
Appendix L Denoising of Preference Graphs for GED
@@
Appendix M Rank ensemble method
@@
Appendix O Prompt Template

### v3

@@ A.5 Definition of Cycle Rate
A.6 Models Used for Ranking
We evaluate model ranking using 30 widely used models from the AlpacaEval dataset, covering diverse architectures and capabilities. Our model set $\mathcal{M}$ includes families such as Llama (Touvron et al., 2023), Vicuna (Zheng et al., 2023), GPT (Achiam et al., 2023), Claude (Anthropic, 2024), Qwen (Yang et al., 2024a), Mistral (Jiang et al., 2023), Yi (Young et al., 2024), and WizardLM (Xu et al., 2023), ensuring a broad assessment. Proprietary models include OpenAI’s GPT series (e.g., gpt-3.5-turbo-0301, gpt-4o-2024-05-13) and Anthropic’s Claude models (e.g., claude-2, claude-3-opus-20240229), serving as strong baselines. Open-source models include multiple versions of Llama (e.g., Meta-Llama-3-8B-Instruct, llama-2-70b-chat-hf) and Qwen (e.g., Qwen1.5-72B-Chat, Qwen2-72B). We also incorporate Mistral models (e.g., Mistral-7B-Instruct-v0.2, Mixtral-8x22B-Instruct-v0.1), reflecting advances in mixture-of-experts architectures. Additional selections include gemini-pro, tulu-2-dpo-70b, oasst-sft-llama-33b, dbrx-instruct, wizardlm-13b, and Yi-34B-Chat, contributing to ranking diversity. This diverse model set enables a comprehensive evaluation of ranking accuracy using AlpacaEval’s ground-truth rankings.
A.7 Dataset
@@
A.8 Aggregation Process in GED Across Different Tasks
@@ Appendix B Alternative Evaluator Configurations for Model Ranking
Appendix C Comparison with Point-Wise Scoring Methods
To further investigate the difference between point-wise and preference-based scoring methods, we conducted additional experiments under the response selection setting. Specifically, we evaluated a point-wise baseline where evaluators assigned scores to responses on a scale of 1 to 5, selecting the highest-rated response. The results are shown in Table 6.
Table 6: Comparison of point-wise methods and GED on response selection tasks.
Model HumanEval AlpacaEval MATH GSM8k GAIA Avg
Point-wise - Llama3-8B 60.97 28.77 74.06 82.65 10.97 51.48
Point-wise - Mistral-7B 62.83 26.93 74.02 83.21 10.22 51.44
Point-wise - Qwen2-7B 60.41 28.30 74.32 84.27 10.82 51.62
Point-wise - Majority Voting 63.72 29.42 74.71 84.36 11.32 52.70
GED (Llama3-8B, Mistral-7B, Qwen2-7B) 70.73 32.44 75.58 88.18 13.33 56.05
From the results, we observe that point-wise methods consistently underperform compared to GED across all tasks. This can be attributed to the fact that point-wise scoring assesses each response independently, without considering the relative quality of different responses. As a result, it lacks the global information available in preference-based ranking, where responses are directly compared. While majority voting improves point-wise performance slightly by aggregating scores from multiple evaluators, it remains inferior to GED, which leverages pairwise preferences to construct a more informed global ranking. Despite its limitations, the point-wise approach has an advantage in terms of computational efficiency, as it requires fewer comparisons than pairwise methods. This trade-off suggests that the choice between point-wise and preference-based scoring should depend on the specific application: point-wise scoring may be preferable in large-scale settings where efficiency is critical, while GED is more effective for achieving higher-quality rankings.
Appendix D Impact of Evaluator Quantity on Denoising Quality
@@
Set of candidate answers $\{ans_{1},ans_{2},\ldots,ans_{n}\}$ for each question $q\in Q$, set of evaluators $\mathcal{A}=\{a_{1},a_{2},\ldots,a_{k}\}$
Set of preference graph sets $\{G_{a}:a\in\mathcal{A}\}$ for each question $q\in Q$
for each question $q\in Q$ do
     for each evaluator $a\in\mathcal{A}$ do
         Initialize vertex set $V_{q}=\{v_{1},v_{2},\ldots,v_{n}\}$, where each $v_{i}$ corresponds to $ans_{i}$
         Initialize edge set $A_{a}=\emptyset$, and weight function $w_{a}:A_{a}\rightarrow\mathbb{R}^{+}$
         for each pair of answers $(ans_{i},ans_{j})$ with $i\neq j$ do
              if $a(ans_{i},ans_{j})>0$ then
                  if $(v_{i},v_{j})\notin A_{a}$ then
                       Add directed edge $(v_{i}\rightarrow v_{j})$ to $A_{a}$, set $w_{a}(v_{i},v_{j})=1$
                  else
                       Increment $w_{a}(v_{i},v_{j})$ by 1
                  end if
              else
                  if $(v_{j},v_{i})\notin A_{a}$ then
                       Add directed edge $(v_{j}\rightarrow v_{i})$ to $A_{a}$, set $w_{a}(v_{j},v_{i})=1$
                  else
                       Increment $w_{a}(v_{j},v_{i})$ by 1
                  end if
              end if
         end for
         Store the preference graph $G_{a}=(V_{q},A_{a},w_{a})$ for evaluator $a$
     end for
end for
Appendix E Construction of the Preference Graph
@@
Appendix F The Importance of Addressing Cyclic Inconsistency
Cyclic inconsistencies in preference data introduce contradictions that undermine meaningful ranking and evaluation. Many applications, including alignment optimization (Ouyang et al., 2022; Song et al., 2024), recommendation systems (Hou et al., 2024), and model evaluation (Liu et al., 2024), rely on transitive preferences for stability and interpretability. For example, OpenAI’s RLHF pipelines (Ouyang et al., 2022) construct training data based on transitive preferences (e.g., A $\succ$ B $\succ$ C), which ensures consistency in model learning. However, if preference cycles exist (e.g., A $\succ$ B $\succ$ C $\succ$ A), the resulting contradictions lead to instability and degrade training performance. In our experiments, removing cycles and transforming preference graphs into directed acyclic graphs (DAGs) improved results across multiple tasks, including response selection, model ranking, and instruction tuning. These improvements were consistent across widely used benchmarks, suggesting that most observed cycles are artifacts of noise rather than meaningful violations of transitivity. Ensuring acyclicity enhances dataset reliability and improves model effectiveness in ranking and evaluation tasks.
Appendix G Proof of Theorem 1
@@
We will then use contradiction to show it is impossible for $\mathrm{MAS}(G)$ to contain neither $(u\rightarrow v)$ nor $(v\rightarrow u)$. Suppose this is true instead. Then, since $\mathrm{MAS}(G)$ is a maximum acyclic subgraph, adding either $(u\rightarrow v)$ or $(v\rightarrow u)$ to $\mathrm{MAS}(G)$ will make it cyclic. That is, $\mathrm{MAS}(G)$ contains a path that goes from $v$ to $u$; meanwhile, it also contains a path from $u$ to $v$. As a result, it contains a cycle that goes from $v$ to $u$ and then from $u$ to $v$, which contradicts with the fact that $\mathrm{MAS}(G)$ is a maximum acyclic subgraph. Therefore, $\mathrm{MAS}(G)$ must contain exactly one of $(u\rightarrow v)$ and $(v\rightarrow u)$. ∎
Appendix H Impact of Evaluator Weighting on GED Performance
@@
Appendix I Expanded Related Work on Preference Denoising for LLMs
@@
This work situates itself within the broader field of denoising preference evaluations for LLMs, addressing inconsistencies and noise in preference graphs. We acknowledge that existing literature has explored two primary approaches to this challenge: within-model preference modeling and modular pre-processing. Below, we provide a detailed comparison of GED with representative methods from these approaches. Within-model approaches, such as robust DPO (rDPO) (Chowdhury et al., 2024) and conservative DPO (cDPO) (Mitchell, 2023), focus on integrating denoising mechanisms directly within the preference modeling process. These methods incorporate regularization techniques to mitigate the effects of noisy or adversarial preference data during model alignment. While these approaches are effective in refining the preference modeling pipeline, they are tightly coupled with specific models and tasks, limiting their versatility. In contrast, GED is a modular framework that operates as a pre-processing step, denoising preference graphs before downstream tasks, making it adaptable to a wide range of applications and models. Modular approaches like CURATRON (Naresh et al., 2024) address noise and missing comparisons in preference datasets using techniques such as low-rank matrix completion. While CURATRON effectively mitigates certain types of noise, it does not explicitly target cyclic inconsistencies (e.g., A $\succ$ B, B $\succ$ C, C $\succ$ A), which are a critical focus of GED. By leveraging a graph-based framework, GED detects and removes such cycles through its denoising process, ensuring that the resulting preference graph is acyclic and thus more consistent and reliable for downstream use. Additionally, GED distinguishes itself by providing theoretical guarantees for recovering the ground truth preference DAG under certain conditions. Furthermore, the ensemble mechanism in GED demonstrates the novel insight that combining small evaluators can surpass the performance of a single stronger evaluator, a feature not emphasized in the aforementioned methods.
Appendix J Impact of Evaluator Selection on GED Performance
@@
The selection of weak evaluators is a critical factor influencing the performance of GED. While our initial experiments demonstrated the benefits of combining weak evaluators, further analysis is necessary to understand how their diversity and capabilities affect GED’s effectiveness. To address this, we conducted additional experiments evaluating three different configurations of evaluator sets: the original set comprising Llama3-8B, Mistral-7B, and Qwen2-7B; a diverse set that adds Gemma-9B [^5] https://huggingface.co/google/gemma-2-9b-it to introduce more variation in model architecture and training data; and a higher-capability set that replaces smaller models with larger ones, specifically Llama3-70B, Mistral-8×7B, and Qwen2-72B. The results, shown in Table 9, reveal two key insights. First, incorporating diversity by adding Gemma-9B improves GED’s performance slightly, increasing the average score from 56.05% to 56.39%. This suggests that models with diverse training data and architectures can contribute to more robust aggregated evaluations. Second, replacing smaller evaluators with higher-capability models yields a more substantial improvement, with the average score rising to 60.89%. Notably, benchmarks like AlpacaEval and GAIA benefit the most from the advanced reasoning and language understanding capabilities of larger models. These findings demonstrate the importance of both diversity and capability in evaluator selection. Diversity provides marginal gains by bringing varied perspectives, while higher-capacity models contribute to more significant improvements by enhancing the overall quality of evaluations. This suggests that, when computational resources allow, incorporating advanced models into the evaluator set can meaningfully boost GED’s performance.
@@
Appendix K Cost Considerations and Active Learning with ActiveGED
Evaluators Set of GED HumanEval AlpacaEval MATH GSM8k GAIA Avg.
Random 57.93 29.58 72.75 84.67 11.52 51.29
ActiveGED (30%) 67.28 30.96 74.65 85.73 11.39 54.00
ActiveGED (50%) 68.62 31.91 74.87 87.06 12.08 54.91
GED (Full Budget) 70.73 32.44 75.58 88.18 13.33 56.05
@@
Appendix L Mitigating Evaluator Biases in GED
@@
Appendix M Denoising of Preference Graphs for GED
@@
Appendix N Rank ensemble method
@@
Appendix P Prompt Template